OpenStudy (anonymous):

A sample of argon gas in a 4.00 L container at 23°C has a mass of 15.98 g. A sample of an unknown gas in the same container at 23°C and at the same pressure has a mass of 8.07 g. What is the molar mass of the unknown gas?

OpenStudy (mos1635):

Argon: P*V=(m1/Mr1)*R*T uknown: P*V=(m2/Mr2)*R*T that give us (m1/Mr1)=(m2/Mr2) Mr2=m2*Mr1/m1 Mr2=.......